# A lower bound for the Weisfeiler-Leman dimension of circulant graphs

### v1

@@ 8.2. A coset S-ring associated with the graph ${\cal G}$
(41) ${\cal A}=({\mathbb{Z}}C_{n_{1}}\wr{\mathbb{Z}}C_{n_{2}}\wr\cdots\wr{\mathbb{Z}}C_{n_{b}})\otimes({\mathbb{Z}}C_{n_{b+1}}\wr{\mathbb{Z}}C_{n_{b+2}}\wr\cdots\wr{\mathbb{Z}}C_{n_{2b}}),$
where the numbers $n_{i}$ are as in Theorem 1.1. For any vertex $v\in V$ of the graph ${\cal G}$, we denote by $G_{v}$ the (cyclic) subgroup of $G$ such that
$|G_{v}|=\begin{cases}n_{1}\cdot\ldots\cdot n_{i}&\text{if $v=u_{i}$ with $1\leq i\leq b$,}\\ n_{b+1}\cdot\ldots\cdot n_{b+j}&\text{if $v=w_{j}$ with $1\leq j\leq b$.}\\ \end{cases}$
Clearly, $G=G_{u_{b}}\times G_{w_{b}}$ and ${\cal A}$ is the tensor product of the iterated wreath products ${\cal A}_{G_{u_{b}}}$ and ${\cal A}_{G_{w_{b}}}$ of the group rings ${\mathbb{Z}}C_{n_{i}}$. The statement below immediately follows from remarks in Subsection 5.9) and Theorem 3.1.
@@
The S-ring ${\cal A}$ is a coset one and the (circulant) scheme ${\cal X}({\cal A})$ is binary.
We enlarge the sets $V_{1}$ and $V_{2}$ by adding to them new elements $u_{0}$ and $w_{0}$, respectively, and set $G_{u_{0}}=G_{w_{0}}=1$. Then $G_{v}$ is an ${\cal A}$-subgroup for each $v$, including these new elements. Moreover,
@@
We sometimes identify the subgroups $G_{u_{i}}$ and $G_{u_{i}}\times G_{w_{0}}$, and $G_{w_{j}}$ with $G_{u_{0}}\times G_{w_{j}}$. Below, for any vertex $v$ of the graph ${\cal G}$, we define the section
${\mathfrak{s}}_{v}=\begin{cases}G_{u_{i}}/G_{u_{i-1}}&\text{if $v=u_{i}$ with $1\leq i\leq b$,}\\ G_{w_{j}}/G_{w_{j-1}}&\text{if $v=w_{j}$ with $1\leq j\leq b$.}\\ \end{cases}$
@@
    Each ${\cal A}$-subgroup is of the form $U\times W$, where $G_{u_{i-1}}\leq U\leq G_{u_{i}}$ and $G_{w_{j-1}}\leq W\leq G_{w_{j}}$ for some $1\leq i,j\leq b$,
@@
    ${\mathfrak{s}}\in{\mathfrak{S}}_{0}({\cal A})$ if and only if ${\mathfrak{s}}\preceq{\mathfrak{s}}_{u,w}$ for some $u\in V_{1}$ and $w\in V_{2}$, where ${\mathfrak{s}}_{u,w}={\mathfrak{s}}_{u}\times{\mathfrak{s}}_{w}$.
@@
Since the group $G=G_{u_{b}}\times G_{w_{b}}$ is cyclic, any subgroup $H\leq G$ is the direct product of two subgroups $U:=H\cap G_{u_{b}}$ and $W:=H\cap G_{w_{b}}$. Assume that $H$ is an ${\cal A}$-subgroup. Since obviously $G_{u_{b}}$ and $G_{w_{b}}$ are also ${\cal A}$-subgroups, so is $U$ and hence it is an ${\cal A}_{G_{u_{b}}}$-subgroup. By the first part of formula (42), there is the smallest $i\in\{1,\ldots,b\}$ such that $U\leq G_{u_{i}}$. Without loss of generality, we assume that $U\neq 1$. Then $U$ is an $({\cal A}_{i-1}\wr{\mathbb{Z}}C_{n_{i}})$-subgroup, where ${\cal A}_{i-1}$ is the restriction of ${\cal A}$ to $G_{u_{i-1}}$. The minimality of $i$ implies that $U\geq G_{u_{i-1}}$ (see Subsection 5.4). Thus, $G_{u_{i-1}}\leq U\leq G_{u_{i}}$. Similarly, one can prove that $G_{w_{j-1}}\leq W\leq G_{w_{j}}$ for some $1\leq j\leq b$. Statement (1) is proved.
To prove statement (2), we note that the necessity is a consequence of the fact that ${\mathfrak{s}}_{u,w}$ is a principal section. Conversely, let ${\mathfrak{s}}$ be an ${\cal A}$-section. By statement (1), there exist subgroups $U^{\prime}\leq U\leq G_{u_{b}}$ and $W^{\prime}\leq W\leq G_{w_{b}}$ such that we have
@@
It follows that ${\mathfrak{s}}={\mathfrak{s}}_{1}\times{\mathfrak{s}}_{2}$, where ${\mathfrak{s}}_{1}=U/U^{\prime}$ and ${\mathfrak{s}}_{2}=W/W^{\prime}$. Assume that ${\mathfrak{s}}\in{\mathfrak{S}}_{0}({\cal A})$. Let us verify that ${\mathfrak{s}}_{1}\preceq{\mathfrak{s}}_{u}$ for some $u\in V_{1}$; the fact that ${\mathfrak{s}}_{2}\preceq{\mathfrak{s}}_{w}$ for some $w\in V_{2}$, is proved similarly.
By [13, Theorem 8.3], we have ${\cal A}_{\mathfrak{s}}={\mathbb{Z}}\,{\mathfrak{s}}$ and hence ${\cal A}_{{\mathfrak{s}}_{1}}={\mathbb{Z}}\,{\mathfrak{s}}_{1}$. Without loss of generality, we may assume that $U\neq 1$ (otherwise one can take $u=1$). Then $G_{u_{i-1}}<U\leq G_{u_{i}}$ for some $1\leq i\leq b$. Now if $U^{\prime}<G_{u_{i-1}}$, then setting ${\mathfrak{t}}=G_{u_{i-1}}/U^{\prime}$ and ${\mathfrak{t}}^{\prime}=U/G_{u_{i-1}}$, we obtain
${\mathbb{Z}}\,{\mathfrak{s}}_{1}={\cal A}_{{\mathfrak{s}}_{1}}={\cal A}_{{\mathfrak{t}}}\wr{\cal A}_{{\mathfrak{t}}^{\prime}}\neq{\mathbb{Z}}\,{\mathfrak{s}}_{1},$
@@
In the following lemma, we show that the inner ${\cal A}$-multipliers are in a one-to-one correspondence with the elements of the direct product of the groups $\operatorname{Aut}({\mathfrak{s}}_{v})$, $v\in V$.
@@
Put ${\mathfrak{S}}=\{{\mathfrak{s}}_{v}:\ v\in V\}$ and ${\mathfrak{M}}=\operatorname{{\mathfrak{M}}}({\cal A})$. Then
@@
The necessity in statement (1) is trivial. To prove the sufficiency, we need to find an inner ${\cal A}$-multiplier ${\mathfrak{m}}$ such that ${\mathfrak{m}}_{\mathfrak{S}}=\bar{m}$. First, put
@@
for all $u\in V_{1}$ and $w\in V_{2}$. Now let section ${\mathfrak{s}}\in{\mathfrak{S}}_{0}({\cal A})$ be arbitrary. By Lemma 8.3(2), there exist $u$ and $w$ such that ${\mathfrak{s}}\preceq{\mathfrak{s}}_{u,w}$. Put
@@
From the above, it follows that the family ${\mathfrak{m}}$ defined by formulas (43) and (44) satisfies the first condition in (17) and the equality $\bar{\mathfrak{m}}={\mathfrak{m}}_{\mathfrak{S}}$. To prove that ${\mathfrak{m}}$ is an inner ${\cal A}$-multiplier, we need to verify the second condition in (17). Let ${\mathfrak{s}}$ and ${\mathfrak{t}}$ be projectively equivalent ${\cal A}$-sections belonging to ${\mathfrak{S}}_{0}({\cal A})$. Then without loss of generality we may assume that both ${\mathfrak{s}}$ and ${\mathfrak{t}}$ are subsections of a section ${\mathfrak{s}}_{u,w}$ for some $u$ and $w$. The automorphism ${\mathfrak{m}}_{{\mathfrak{s}}_{u,w}}$ of the cyclic group ${\mathfrak{s}}_{u,w}$ acts by raising to a certain integer power $\lambda$. Hence the automorphisms ${\mathfrak{m}}_{\mathfrak{s}}=({\mathfrak{m}}_{{\mathfrak{s}}_{u,w}})^{\mathfrak{s}}$ and ${\mathfrak{m}}_{\mathfrak{t}}=({\mathfrak{m}}_{{\mathfrak{s}}_{u,w}})^{\mathfrak{t}}$ act by raising to the powers $\lambda\pmod{|{\mathfrak{s}}|}$ and $\lambda\pmod{|{\mathfrak{t}}|}$, respectively. Since $|{\mathfrak{s}}|=|{\mathfrak{t}}|$ and the isomorphism $f_{{\mathfrak{s}},{\mathfrak{t}}}:{\mathfrak{s}}\to{\mathfrak{t}}$ preserves raising to a power, we conclude that $({\mathfrak{m}}_{\mathfrak{s}})^{f_{{\mathfrak{s}},{\mathfrak{t}}}}={\mathfrak{m}}_{\mathfrak{t}}$, as required.
To prove statement (2), assume that ${\mathfrak{m}}\in{\mathfrak{M}}$ is such that ${\mathfrak{m}}_{\mathfrak{s}}=\operatorname{id}_{\mathfrak{s}}$ for all ${\mathfrak{s}}\in{\mathfrak{S}}$. Note that if $u\in V_{1}$ and $w\in V_{2}$, then ${\mathfrak{m}}_{{\mathfrak{s}}_{u,w}}=({\mathfrak{m}}_{{\mathfrak{s}}_{u}},{\mathfrak{m}}_{{\mathfrak{s}}_{w}})=\operatorname{id}_{{\mathfrak{s}}_{u,w}}$. Now let ${\mathfrak{s}}\in{\mathfrak{S}}_{0}({\cal A})$. By Lemma 8.3(2), we have ${\mathfrak{s}}\preceq{\mathfrak{s}}_{u,w}$ for some $u$ and $w$. By the first condition in (17), this implies that
@@
Thus the ${\mathfrak{m}}$ is the trivial inner ${\cal A}$-multiplier and we are done. ∎
The basic sets of the S-ring ${\cal A}$ can conveniently be divided into parts parameterized by the ${\cal A}$-sets defined below; this parameterization will be used in Section 10. Namely, by Lemma 8.3, for any two indices $1\leq i,j\leq b$, there is an ${\cal A}$-subgroup $G_{i,j}=G_{u_{i}}\times G_{w_{j}}$. It follows that the difference
@@
is an ${\cal A}$-set. Since any nonidentity element of $G$ belongs to some of these differences, the union of all of them is equal to $G^{\#}$ and each nonidentity set $X\in{\cal S}({\cal A})$ is contained in at least one of them. Moreover, it is not hard to verify that
@@
8.3. Inner ${\cal A}$-multipliers associated with ${\cal G}$
Let $v=v_{i}$ be a vertex of the graph ${\cal G}$, $1\leq i\leq a$. Then ${\mathfrak{s}}_{v}=C_{n_{i}}$ is a (cyclic) group of order $n_{i}=p_{i}q_{i}$, where $p_{i}$ and $q_{i}$ are distinct odd primes, $p_{i}<q_{i}$. It follows that the group
@@ 8.3. Inner ${\cal A}$-multipliers associated with ${\cal G}$
Let ${\mathfrak{M}}=\operatorname{{\mathfrak{M}}}({\cal A})$ and $f\in K_{V}$. Then there is a unique inner ${\cal A}$-multiplier ${\mathfrak{m}}_{f}\in{\mathfrak{M}}$ such that $({\mathfrak{m}}_{f})_{{\mathfrak{s}}_{v}}=\pi_{v}(f_{v})$ for all $v\in V$, Moreover, the mapping

### v2

@@ 8.2. A coset S-ring associated with the graph $\mathcal{G}$}
(41) ${\mathcal{A}}=({\mathbb{Z}}C_{n_{1}}\wr{\mathbb{Z}}C_{n_{2}}\wr\cdots\wr{\mathbb{Z}}C_{n_{b}})\otimes({\mathbb{Z}}C_{n_{b+1}}\wr{\mathbb{Z}}C_{n_{b+2}}\wr\cdots\wr{\mathbb{Z}}C_{n_{2b}}),$
where the numbers $n_{i}$ are as in Theorem 1.1. For any vertex $v\in V$ of the graph $\mathcal{G}$}, we denote by $G_{v}$ the (cyclic) subgroup of $G$ such that
[…]
Clearly, $G=G_{u_{b}}\times G_{w_{b}}$ and $\mathcal{A}$} is the tensor product of the iterated wreath products ${\mathcal{A}}_{G_{u_{b}}}$ and ${\mathcal{A}}_{G_{w_{b}}}$ of the group rings ${\mathbb{Z}}C_{n_{i}}$. The statement below immediately follows from remarks in Subsection 5.9) and Theorem 3.1.
@@
The S-ring $\mathcal{A}$} is a coset one and the (circulant) scheme ${\mathcal{X}}({\mathcal{A}})$ is binary.
We enlarge the sets $V_{1}$ and $V_{2}$ by adding to them new elements $u_{0}$ and $w_{0}$, respectively, and set $G_{u_{0}}=G_{w_{0}}=1$. Then $G_{v}$ is an $\mathcal{A}$}-subgroup for each $v$, including these new elements. Moreover,
@@
We sometimes identify the subgroups $G_{u_{i}}$ and $G_{u_{i}}\times G_{w_{0}}$, and $G_{w_{j}}$ with $G_{u_{0}}\times G_{w_{j}}$. Below, for any vertex $v$ of the graph $\mathcal{G}$}, we define the section
[…]
@@
    Each $\mathcal{A}$}-subgroup is of the form $U\times W$, where $G_{u_{i-1}}\leq U\leq G_{u_{i}}$ and $G_{w_{j-1}}\leq W\leq G_{w_{j}}$ for some $1\leq i,j\leq b$,
@@
    ${\mathfrak{s}}\in{\mathfrak{S}}_{0}({\mathcal{A}})$ if and only if ${\mathfrak{s}}\preceq{\mathfrak{s}}_{u,w}$ for some $u\in V_{1}$ and $w\in V_{2}$, where ${\mathfrak{s}}_{u,w}={\mathfrak{s}}_{u}\times{\mathfrak{s}}_{w}$.
@@
Since the group $G=G_{u_{b}}\times G_{w_{b}}$ is cyclic, any subgroup $H\leq G$ is the direct product of two subgroups $U:=H\cap G_{u_{b}}$ and $W:=H\cap G_{w_{b}}$. Assume that $H$ is an $\mathcal{A}$}-subgroup. Since obviously $G_{u_{b}}$ and $G_{w_{b}}$ are also $\mathcal{A}$}-subgroups, so is $U$ and hence it is an ${\mathcal{A}}_{G_{u_{b}}}$-subgroup. By the first part of formula (42), there is the smallest $i\in\{1,\ldots,b\}$ such that $U\leq G_{u_{i}}$. Without loss of generality, we assume that $U\neq 1$. Then $U$ is an $({\mathcal{A}}_{i-1}\wr{\mathbb{Z}}C_{n_{i}})$-subgroup, where ${\mathcal{A}}_{i-1}$ is the restriction of $\mathcal{A}$} to $G_{u_{i-1}}$. The minimality of $i$ implies that $U\geq G_{u_{i-1}}$ (see Subsection 5.4). Thus, $G_{u_{i-1}}\leq U\leq G_{u_{i}}$. Similarly, one can prove that $G_{w_{j-1}}\leq W\leq G_{w_{j}}$ for some $1\leq j\leq b$. Statement (1) is proved.
To prove statement (2), we note that the necessity is a consequence of the fact that ${\mathfrak{s}}_{u,w}$ is a principal section. Conversely, let $\mathfrak{s}$} be an $\mathcal{A}$}-section. By statement (1), there exist subgroups $U^{\prime}\leq U\leq G_{u_{b}}$ and $W^{\prime}\leq W\leq G_{w_{b}}$ such that we have
@@
It follows that ${\mathfrak{s}}={\mathfrak{s}}_{1}\times{\mathfrak{s}}_{2}$, where ${\mathfrak{s}}_{1}=U/U^{\prime}$ and ${\mathfrak{s}}_{2}=W/W^{\prime}$. Assume that ${\mathfrak{s}}\in{\mathfrak{S}}_{0}({\mathcal{A}})$. Let us verify that ${\mathfrak{s}}_{1}\preceq{\mathfrak{s}}_{u}$ for some $u\in V_{1}$; the fact that ${\mathfrak{s}}_{2}\preceq{\mathfrak{s}}_{w}$ for some $w\in V_{2}$, is proved similarly.
By [13, Theorem 8.3], we have ${\mathcal{A}}_{\mathfrak{s}}={\mathbb{Z}}\,{\mathfrak{s}}$ and hence ${\mathcal{A}}_{{\mathfrak{s}}_{1}}={\mathbb{Z}}\,{\mathfrak{s}}_{1}$. Without loss of generality, we may assume that $U\neq 1$ (otherwise one can take $u=1$). Then $G_{u_{i-1}}<U\leq G_{u_{i}}$ for some $1\leq i\leq b$. Now if $U^{\prime}<G_{u_{i-1}}$, then setting ${\mathfrak{t}}=G_{u_{i-1}}/U^{\prime}$ and ${\mathfrak{t}}^{\prime}=U/G_{u_{i-1}}$, we obtain
${\mathbb{Z}}\,{\mathfrak{s}}_{1}={\mathcal{A}}_{{\mathfrak{s}}_{1}}={\mathcal{A}}_{{\mathfrak{t}}}\wr{\mathcal{A}}_{{\mathfrak{t}}^{\prime}}\neq{\mathbb{Z}}\,{\mathfrak{s}}_{1},$
@@
In the following lemma, we show that the inner $\mathcal{A}$}-multipliers are in a one-to-one correspondence with the elements of the direct product of the groups $\operatorname{Aut}({\mathfrak{s}}_{v})$, $v\in V$.
@@
Put ${\mathfrak{S}}=\{{\mathfrak{s}}_{v}:\ v\in V\}$ and ${\mathfrak{M}}=\operatorname{{\mathfrak{M}}}({\mathcal{A}})$. Then
@@
The necessity in statement (1) is trivial. To prove the sufficiency, we need to find an inner $\mathcal{A}$}-multiplier $\mathfrak{m}$} such that ${\mathfrak{m}}_{\mathfrak{S}}=\bar{m}$. First, put
@@
for all $u\in V_{1}$ and $w\in V_{2}$. Now let section ${\mathfrak{s}}\in{\mathfrak{S}}_{0}({\mathcal{A}})$ be arbitrary. By Lemma 8.3(2), there exist $u$ and $w$ such that ${\mathfrak{s}}\preceq{\mathfrak{s}}_{u,w}$. Put
@@
From the above, it follows that the family $\mathfrak{m}$} defined by formulas (43) and (44) satisfies the first condition in (17) and the equality $\bar{\mathfrak{m}}={\mathfrak{m}}_{\mathfrak{S}}$. To prove that $\mathfrak{m}$} is an inner $\mathcal{A}$}-multiplier, we need to verify the second condition in (17). Let $\mathfrak{s}$} and $\mathfrak{t}$} be projectively equivalent $\mathcal{A}$}-sections belonging to ${\mathfrak{S}}_{0}({\mathcal{A}})$. Then without loss of generality we may assume that both $\mathfrak{s}$} and $\mathfrak{t}$} are subsections of a section ${\mathfrak{s}}_{u,w}$ for some $u$ and $w$. The automorphism ${\mathfrak{m}}_{{\mathfrak{s}}_{u,w}}$ of the cyclic group ${\mathfrak{s}}_{u,w}$ acts by raising to a certain integer power $\lambda$. Hence the automorphisms ${\mathfrak{m}}_{\mathfrak{s}}=({\mathfrak{m}}_{{\mathfrak{s}}_{u,w}})^{\mathfrak{s}}$ and ${\mathfrak{m}}_{\mathfrak{t}}=({\mathfrak{m}}_{{\mathfrak{s}}_{u,w}})^{\mathfrak{t}}$ act by raising to the powers $\lambda\pmod{|{\mathfrak{s}}|}$ and $\lambda\pmod{|{\mathfrak{t}}|}$, respectively. Since $|{\mathfrak{s}}|=|{\mathfrak{t}}|$ and the isomorphism $f_{{\mathfrak{s}},{\mathfrak{t}}}:{\mathfrak{s}}\to{\mathfrak{t}}$ preserves raising to a power, we conclude that $({\mathfrak{m}}_{\mathfrak{s}})^{f_{{\mathfrak{s}},{\mathfrak{t}}}}={\mathfrak{m}}_{\mathfrak{t}}$, as required.
To prove statement (2), assume that ${\mathfrak{m}}\in{\mathfrak{M}}$ is such that ${\mathfrak{m}}_{\mathfrak{s}}=\operatorname{id}_{\mathfrak{s}}$ for all ${\mathfrak{s}}\in{\mathfrak{S}}$. Note that if $u\in V_{1}$ and $w\in V_{2}$, then ${\mathfrak{m}}_{{\mathfrak{s}}_{u,w}}=({\mathfrak{m}}_{{\mathfrak{s}}_{u}},{\mathfrak{m}}_{{\mathfrak{s}}_{w}})=\operatorname{id}_{{\mathfrak{s}}_{u,w}}$. Now let ${\mathfrak{s}}\in{\mathfrak{S}}_{0}({\mathcal{A}})$. By Lemma 8.3(2), we have ${\mathfrak{s}}\preceq{\mathfrak{s}}_{u,w}$ for some $u$ and $w$. By the first condition in (17), this implies that
@@
Thus the $\mathfrak{m}$} is the trivial inner $\mathcal{A}$}-multiplier and we are done. ∎
The basic sets of the S-ring $\mathcal{A}$} can conveniently be divided into parts parameterized by the $\mathcal{A}$}-sets defined below; this parameterization will be used in Section 10. Namely, by Lemma 8.3, for any two indices $1\leq i,j\leq b$, there is an $\mathcal{A}$}-subgroup $G_{i,j}=G_{u_{i}}\times G_{w_{j}}$. It follows that the difference
@@
is an $\mathcal{A}$}-set. Since any nonidentity element of $G$ belongs to some of these differences, the union of all of them is equal to $G^{\#}$ and each nonidentity set $X\in{\mathcal{S}}({\mathcal{A}})$ is contained in at least one of them. Moreover, it is not hard to verify that
@@
8.3. Inner $\mathcal{A}$}-multipliers associated with $\mathcal{G}$}
Let $v=v_{i}$ be a vertex of the graph $\mathcal{G}$}, $1\leq i\leq a$. Then ${\mathfrak{s}}_{v}=C_{n_{i}}$ is a (cyclic) group of order $n_{i}=p_{i}q_{i}$, where $p_{i}$ and $q_{i}$ are distinct odd primes, $p_{i}<q_{i}$. It follows that the group
@@ 8.3. Inner $\mathcal{A}$}-multipliers associated with $\mathcal{G}$}
Let ${\mathfrak{M}}=\operatorname{{\mathfrak{M}}}({\mathcal{A}})$ and $f\in K_{V}$. Then there is a unique inner $\mathcal{A}$}-multiplier ${\mathfrak{m}}_{f}\in{\mathfrak{M}}$ such that $({\mathfrak{m}}_{f})_{{\mathfrak{s}}_{v}}=\pi_{v}(f_{v})$ for all $v\in V$, Moreover, the mapping